The value of ∆G\deg for formation of Cr_(2)O_(3) is -540 kJmol^(-1) and that of Al_(2)O_(3) is -827 kJmol^(-1). Is the reduction of Cr_(2)O3 possible with Al ?
1
Expert's answer
2020-08-31T07:04:51-0400

Yes, the reduction of Cr_(2)_O_(3) with Al is possible. ∆G ( of this reaction) <0. ∆G=-827-(-540)= -827+540=-287